OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = 3x + 1 A. (6, 18) and (0, 1) B. (6, 19) and (0, 0) C. (19, 6) and (0, 1) D. (6, 19) and (0, 1)

sammixboo (sammixboo):

Plug them in :) Choice A. 18 = 3*6 + 1 1 = 3*0 + 1 Choice B 19 = 3*6 + 1 0 = 3*0 + 1 Choice C 6 = 8*19 + 1 1 = 3*0 + 1 Choice D 19 = 3*6 + 1 1 = 3*0 + 1

OpenStudy (anonymous):

Substitute your (x, y) into the equation.

OpenStudy (anonymous):

okay

OpenStudy (lizz123):

do you get it?

OpenStudy (lizz123):

@1DISAMAZING2345

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@lizz123

OpenStudy (triciaal):

do you still need help?